(x+15).=10 y. where y=5
Answer by unlockmath(1688)

Hello,
For this [(x+15)=10 y] let's plug in the value of y as
(x+15)=10(4) Rewritten as:
(x+15)=40
Subtract 15 from both sides to get:
x=25
Make sense?
RJ
www.math-unlock.com